# DevBench: A Comprehensive Benchmark for Software Development

### v2

@@ 6.1 Setup
[…]
@@
[…]
6.2 Main Results
Results across prompting methods
We first conduct experiments to examine the effects of different prompting methods on a subset of DevBench using three representative models: GPT-4-Turbo, CodeLlama-34B-Instruct, and DeepSeek-Coder-33B-Instruct.
@@ 6.2 Main Results
Table 3 illustrates the results of the implementation task for our study. In general, Execution-Feedback leads to the optimal performance, where GPT-4-Turbo benefits the most, especially on acceptance tests. In contrast, CodeLlama-34B-Instruct and DeepSeek-Coder-33B exhibited no improvement with any review process. We note that the efficacy of the review process could be understated as our automated testing is too rigorous and sparse to reflect the improvements. We observe that, despite no substantial improvements on reference tests, the code quality notably improved with Execution-Feedback prompt. However, there is no significant improvements using the Normal-Review setting compared with the No-Review setting. Models consistently provide encouraging feedback on generated code and offer unhelpful suggestions, such as the addition of unnecessary error handling or reorganization. This indicates that the models are unable to comprehend complicated code by merely reading it, lacking external knowledge or assistance like execution feedback.
@@
For the testing tasks, which are relatively easier than the implementation, there are various observations. In Table 5, we find no clear evidence that the review process brings stable benefits to unit testing. However, on the acceptance testing, Normal-Review brings enhancement to the performance. Regarding the visibility of implementation source code, it is common practice not to expose source code and execution feedback for acceptance testing, while it’s allowed to employ the source code as input for unit testing. As shown in Table 4, models can barely generate executable acceptance testing code and incorporating source code as additional input dramatically increases the performance.
Results across base models
@@
Table 6 illustrates the main results on DevBench with optimal prompting methods applied for each task.555 Execution-Feedback is used for Environment Setup and Implementation; Normal-Review w/ Src Code for Acceptance Testing; No-Review w/ Src Code for Unit Testing as Normal-Review w/ Src Code shows no clear advantage. We find that GPT-4-Turbo demonstrates superior performance compared to other models, while all models evaluated are far from satisfactory. DevBench can effectively distinguish between models of varying capabilities. Smaller models, such as CodeLlama-7B/13B-Instruct and DeepSeek-Coder-1.3B-Instruct, demonstrate inherent limitations, frequently unable to generate syntactically accurate code or follow the instructions. These models tend to generate mere code skeletons or fill the function body with only comments. Larger open-sourced models and GPT models, while generating more reasonable code, still struggle with the subtleties of complex code structure and logic, such as variable type conversion, function arguments and object-oriented class.
@@
Generally, the models’ performances on the implementation task are all below the 10% pass rate. The highest-performing model, GPT-4-Turbo, registers only a 7.1% pass rate on reference acceptance tests and 8.0% on unit tests, while some other models score zero, failing all reference tests. Despite prior research such as HumanEval (Chen et al., 2021) indicating that models can manage simple code-writing tasks, substantial challenges remain in more complex coding scenarios within DevBench. We break down the implementation results into different languages in Figure 5 and find that models particularly struggle in handling Java and C/C++, whose stringent syntax requirements tend to magnify the models’ deficiencies in managing intricate details. This points to the necessity for more enriched and diverse training data across programming languages to bridge this competency gap.
Compared with the implementation task, other tasks are relatively simple but still challenging. Regarding the environment setup task, we note that the test cases for the environment setup task are quite scarce compared to other tasks666 Except for C/C++ and easy Python repositories that are absent, only 12 repositories are involved in the environment setup task. We will resolve this issue in future work. However, DevBench contains rich test cases for other tasks. Table 2 evidences that DevBench features fruitful tests for the implementation task. For the testing generation tasks, our prompts depict fine-grained requirements and ensure the quantity of generated testing cases. , therefore the results are more influenced by the randomness. Roughly speaking, GPT-4-Turbo reaches a 41.7% pass rate in building environments, while open-sourced models largely fall behind it. For the testing tasks, GPT-4-Turbo still obtains the highest scores and open-sourced models perform worse. We identify an outlier that DeepSeek-Coder-6.7B-Instruct achieves the highest score among open-sourced models and approach GPT-4-Turbo in acceptance testing. The model somehow cheats on this task by generating meaningless but executable testing code (see Appendix C.4). With respect to the unit testing, the overall Oracle Test and Coverage scores are quite low, which are averaged across all repositories. However, for those generated testing codes that can be successfully executed (obtain non-zero Oracle Test score), the coverage scores are relatively high, suggesting the models’ promising capability on this problem.
@@
Table 7 shows the results of software design using GPT-4-Turbo as the Judge and GPT-3.5-Turbo as the baseline reference. GPT-4-Turbo dominantly outperforms GPT-3.5-Turbo with extraordinarily high win rates in all cases. Regarding the open-sourced models, as the size increases, models consistently produce higher quality design documents on both metrics, while they are relatively inferior on faithfulness.
We compare the LLM-as-a-Judge results with human majority annotations. Low agreements are observed with tie considered, which aligns with the findings in previous studies (Zheng et al., 2023a). It is reasonable as a tie is hard to define and judge, especially for highly complicated and structured software design documents. Without tie, GPT-4-Turbo reaches 79.2% and 83.2% agreements on the general principles and faithfulness metrics, respectively. This means GPT-4-Turbo’s judgments align with the majority of humans and could serve as a good alternative for automated software design evaluation.
@@
Table 7: Win rate of pairwise comparison against GPT-3.5-Turbo on Software Desgin on a subset of DevBench where results are averaged across different repositories and sub-tasks uniformly. $\dagger$}: the general principles metric. $\ddagger$}: the faithfulness metric. w/ Tie: inconsistent results are considered as a tie. We also report agreement with Human Majority.
@@ 6.3 Analysis
Models struggle with Makefile and Gradle.. LLMs often face challenges in generating accurate Makefile for C/C++ and Gradle files for Java. Frequently, the generated files are deficient in critical components like source code files, necessary dependencies and essential tasks. In C/C++ and Java repositories, approximately 90% of compilation and execution errors can be attributed to these issues. We find that even GPT-4-Turbo occasionally fail in basic syntax errors on compilation files. This is potentially caused by insufficient training data related to these compilation tools.
Models fail to configure function arguments. We observe that models have difficulties in accurately following complex requirements, often failing to correctly configure and utilize function arguments. This contrasts with previous studies that highlight the models’ relative proficiency in implementing function bodies. Such errors in arguments are detected in 20% of the repositories examined using GPT-4-Turbo, in 30% with CodeLlama-34B-Instruct and in 20% with DeepSeek-Coder-33B-Instruct. These issues cannot be effectively corrected through our review process with either Normal-Review or Execution-Feedback settings. This is because models frequently misattribute execution errors to incorrect usage of test programs, despite the clear prompt that the gold testing has already been given.
Models face obstacles in advanced programming skills.. Another observation reveals that models encounter difficulties with advanced programming skills such as overloading, utilizing pointers, and object-oriented programming. In 20% of the repositories tested, there is a noticeable misunderstanding of such advanced notions. Models are unable to distinguish between private and public members/methods, often erroneously attempting to access private variables or import class functions into other code files. Despite receiving detailed feedback on execution errors, models still fail to correctly address these issues, indicating a knowledge gap in complex programming features and skills.
@@ 7 Conclusion
The DevBench framework introduced in this paper presents a leap forward in evaluating LLMs within the domain of automated software development. By employing a multi-stage benchmarking process, DevBench comprehensively assesses LLMs across a spectrum of tasks including design, environment setup, implementation, and testing. Empirical findings reveal that current pre-trained models like GPT-4-Turbo are still confronted with substantial challenges within DevBench. Through analysis, we identify models’ limitations in understanding the complex repository structures and handling the nuanced demands of comprehensive software development. These insights obtained from the DevBench assessments elucidate critical pathways for future code model development.
Impact Statement
DevBench and prior works demonstrate the exciting potential and impact that LLM oriented methods can have towards automating and assisting software engineering. As more automatic prototypes and processes will likely emerge to address different parts of various software engineering workflows, it becomes increasingly important to understand models’ performance on not only each individual task, but also how model output on one task affects subsequent downstream tasks that then receive the generation as input. Especially as LLM powered frameworks are deployed in real-world scenarios, robust and realistic model evaluation is critical to answering whether their designs and implementations are ready to be used and affect human end users. Our work’s novel, holistic evaluation of software engineering incorporates tasks is an important first step towards answering such questions, and we are hopeful that DevBench may inspire more work that covers additional areas of software engineering where LLMs could play a role.
In addition to the task formulations and experimental findings presented in the main paper, we also provide detailed descriptions, motivations, examples, data, and additional results in the Appendix to supplement our claims and demonstrate how other practitioners can easily create their own code repositories and system design documents for evaluation using DevBench. While DevBench’s current 22 repositories provide strong coverage of popular programming languages and use cases, we are eager to continually expand this coverage over time.
@@ Appendix A Software Engineering Tasks with a Running Example
We describe concepts of software engineering tasks using one of the subjects of DevBench, named Actor Relationship Game. The Actor Relationship Game is a Java-based application that allows users to explore connections between popular actors through their movie collaborations, using data from The Movie Database (TMDB) API. It constructs an actor graph and identifies the shortest path of relationships between any two actors.
@@ A.1.3 Architecture Design
.1 /. .2 .gitignore. .2 src. .3 acceptanceTest. .4 java. .5 Actor_relationship_game. .6 ActorRelationshipAcceptanceTest.java. .3 main. .4 java. .5 Actor_relationship_game. .6 Actor.java. .6 Movie.java. .6 ActorGraph.java. .6 ActorGraphUtil.java. .6 GameplayInterface.java. .6 GraphCreation.java. .6 TMDBApi.java. .3 test. .4 java. .5 Actor_relationship_game. .6 ActorTest.java. .6 MovieTest.java. .6 ActorGraphTest.java. .6 ActorGraphUtilTest.java. .6 GameplayInterfaceTest.java. .6 GraphCreationTest.java. .6 TMDBApiTest.java. .2 build.gradle. .2 README.md.
@@ A.3 Quality Assurance
Code Listing LABEL:lst:ut is part of the unit test suite for the Actor Relationship Game Repository, specifically designed to validate the functionality of the Actor class. Using the JUnit framework, it defines two test cases: testActorIdAndName and testMovieIds. The first test, testActorIdAndName, instantiates an Actor object with a specific ID and name (”101” and ”John Doe”, respectively) and asserts that the getId() and getName() methods correctly return these values, ensuring the actor’s identity is accurately stored and retrievable. The second test, testMovieIds, creates another Actor object and adds two movie IDs (”201” and ”202”) to the actor’s list of movie IDs. It then verifies that these movie IDs are indeed associated with the actor by checking if the actor’s getMovieIds() set contains the added IDs. Together, these tests check the integrity of the Actor class’s basic functionalities: maintaining an actor’s identity and managing their associated movie IDs.
@@
In Code Listing LABEL:lst:at, the example acceptance test is designed to verify the functionality of generating and comparing actor lists from graph data. It employs the runGradleTask method to execute specific Gradle tasks for creating graph data and generating actor lists into specified file paths, using parameters for file names to differentiate between reference and test data. The test first runs the runGraphCreation task with paths for both reference and test graphs, followed by the runActorGraphUtil task to generate actor lists from these graphs into specified file paths. Once the actor lists are generated, the test reads lines from both the reference and test actor list files and then iterates through each line in the reference actor list, followed by an assertion that each actor from the reference list is also present in the test list. This process effectively checks the integrity and consistency of the actor list generation feature by ensuring that the test actor list replicates the reference list accurately, thereby validating the application’s capability to process and output graph-related data correctly.
@@
Appendix B Metrics of Software Design Evaluation
Appendix C Discussions

### v3

@@ 4.2 Main Results
Unit Test
Oracle Test§ Oracle Test§ Coverage\$
@@
Table 6: Tasks 2 to 5 results on DevEval.Italic figures: test cases for the Environment Setup task are quite scarce compared to other tasks, therefore the results are more influenced by the randomness6. $\S$§: all results are averaged across all repositories and weighted uniformly. $\P$: all results are averaged across all repositories and weighted by the number of code lines. \$currency-dollar\\$\$: the results on the left side are averaged across all repositories and weighted uniformly, showing the overall scores. The results on the right side in the parenthesis are averaged across all valid repositories and weighted uniformly, where models have generated executable testing code. $\heartsuit$: the model has generated meaningless but executable testing code.
@@
Table 3 illustrates the results of the implementation task for our study. In general, Execution-Feedback leads to the optimal performance, where GPT-4-Turbo benefits the most, especially on acceptance tests. In contrast, CodeLlama-34B-Instruct and DeepSeek-Coder-33B exhibited no improvement with any review process. We note that the efficacy of the review process could be understated as our automated testing is too rigorous and sparse to reflect the improvements. We observe that, despite no substantial improvements on reference tests, the code quality notably improved with Execution-Feedback prompt. However, there is no significant improvements using the Normal-Review setting compared with the No-Review setting. Models consistently provide unhelpful suggestions, such as the addition of unnecessary error handling or reorganization. This indicates that the models are unable to comprehend complicated code by merely reading it, lacking external knowledge like execution feedback.
@@
For the testing tasks, which are relatively easier than the implementation, there are various observations. In Table 4, we find no clear evidence that the review process brings stable benefits to unit testing. However, on the acceptance testing in Table 5, Normal-Review brings enhancement to the performance. The degradation in unit testing performance with review mainly stems from extended input length challenging models’ long-context comprehension. Imprecise reviewer suggestions may also reduce output quality. Regarding the visibility of implementation source code, it is common practice not to expose source code and execution feedback for acceptance testing, while it’s allowed to employ the source code as input for unit testing. As shown in Table 5, models can barely generate executable acceptance testing code and incorporating source code as additional input dramatically increases the performance.
Results across models
@@
Table 6 illustrates the main results on DevEval with optimal prompting methods applied for each task.555 Execution-Feedback is used for Environment Setup and Implementation; Normal-Review w/ Src Code for Acceptance Testing; No-Review w/ Src Code for Unit Testing as Normal-Review w/ Src Code shows no clear advantage. We find that GPT-4-Turbo demonstrates superior performance compared to other models, while all models are far from satisfactory. DevEval can effectively distinguish between models of varying capabilities. Smaller models, such as CodeLlama-7B/13B-Instruct and DeepSeek-Coder-1.3B-Instruct, demonstrate inherent limitations, frequently unable to generate syntactically accurate code or follow the instructions. These models tend to generate mere code skeletons or fill the function body with only comments. Larger open-sourced models and GPT models, while generating more reasonable code, still struggle with the subtleties of complex code structure and logic, such as variable type conversion, function arguments and object-oriented classes.
@@
Generally, the models’ performances on the implementation task are all below the 10% pass rate. The highest-performing model, GPT-4-Turbo, registers only a 7.1% pass rate on reference acceptance tests and 8.0% on unit tests, while some other models score zero, failing all reference tests. Despite prior research such as HumanEval (Chen et al., 2021) indicating that models can manage simple code-writing tasks, substantial challenges remain in more complex coding scenarios within DevEval. We break down the implementation results into different languages in Figure 8 and find that models particularly struggle in handling Java and C/C++, whose stringent syntax requirements tend to magnify the models’ deficiencies in managing intricate details. This points to the necessity for more enriched and diverse training data across programming languages to bridge this competency gap.
Compared with the implementation task, other tasks are relatively simple but still challenging. Regarding the environment setup task, we note that the test cases for the environment setup task are quite scarce compared to other tasks666 Except for C/C++ and easy Python repositories that are absent, only 12 repositories are involved in the environment setup task. We will resolve this issue in future work. However, DevEval contains rich test cases for other tasks. Table 2 evidences that DevEval features fruitful tests for the implementation task. For the testing generation tasks, our prompts depict fine-grained requirements and ensure the quantity of generated testing cases. , therefore the results are more influenced by the randomness. Roughly speaking, GPT-4-Turbo reaches a 41.7% pass rate in building environments, while open-sourced models largely fall behind it. For the testing tasks, GPT-4-Turbo still obtains the highest scores and open-sourced models perform worse. We identify an outlier that DeepSeek-Coder-6.7B-Instruct achieves the highest score among open-sourced models and approach GPT-4-Turbo in acceptance testing. The model somehow cheats on this task by generating meaningless but executable testing code (detail in Section E.4). With respect to the unit testing, the overall Oracle Test and Coverage scores are quite low, which are averaged across all repositories. However, for those generated testing codes that can be successfully executed (obtain non-zero Oracle Test score), the coverage scores are relatively high, suggesting the models’ promising capability on this problem.
@@
Table 7 shows the results of software design using GPT-4-Turbo as the Judge and GPT-3.5-Turbo as the baseline for comparison. More details on how we evaluate the software design are found in Appendix D. GPT-4-Turbo dominantly outperforms GPT-3.5-Turbo with extraordinarily high win rates in all cases. Regarding the open-sourced models, as the size increases, models consistently produce higher quality design documents on both metrics, while they are relatively inferior on faithfulness.
We compare the LLM-as-a-Judge results with human majority annotations. Low agreements are observed with tie considered, which aligns with previous studies (Zheng et al., 2023a). It is reasonable as a tie is hard to define and judge, especially for highly complicated and structured software design documents. Without tie, GPT-4-Turbo reaches 79.2% and 83.2% agreements on the general principles and faithfulness metrics, respectively. This means GPT-4-Turbo’s judgments align with the majority of humans and could serve as a good alternative for automated software design evaluation.
@@ G F G F
Table 7: Win rate of pairwise comparison against GPT-3.5-Turbo on Software Desgin on a subset of DevEval . Results are averaged across different repositories and sub-tasks uniformly. $\dagger$}: general principles. $\ddagger$}: faithfulness. w/ Tie: inconsistent results are considered as a tie. We also report agreement with Human Majority.
@@ 4.3 Analysis
Our experiments reveal several challenges faced by LLMs, such as difficulties in generating accurate Makefiles and Gradle files, handling function redefinitions, and managing file references in multi-file repositories. Models also struggle with correct function parameter usage, naming conventions, type handling, and managing variable scope. Furthermore, they frequently fabricate variables or misinterpret data files, leading to hallucination issues. For more detailed experimental findings and analysis, please refer to the Appendix E.
@@ 5 Conclusion
The DevEval framework presents a leap forward in studying LLMs within the domain of automated software development. By employing a multi-stage evaluation process, DevEval comprehensively assesses LLMs across a spectrum of tasks including design, environment setup, implementation, and testing. Empirical findings reveal that pre-trained models like GPT-4-Turbo are still confronted with substantial challenges within DevEval. Through analysis, we identify models’ limitations in understanding the complex repository structures and handling the nuanced demands of comprehensive software development. These insights elucidate critical pathways for future model development.
6 Limitations
One limitation of our work is the limited number of repositories used in the study, with only 22 curated examples across four programming languages. This relatively small dataset may not fully capture the wide variety of challenges and complexities present in real-world software development. Although we selected diverse repositories to represent different domains and programming paradigms, a broader and more extensive collection of repositories would provide a more comprehensive evaluation of LLM performance. Future work could address this by incorporating additional repositories to better generalize the findings.
Appendix Table of Contents
[…]
[appendices]0
Appendix
@@ Appendix A Software Engineering Tasks with a Running Example
We describe concepts of software engineering tasks using one of the subjects of DevEval, named Actor Relationship Game. The Actor Relationship Game is a Java-based application that allows users to explore connections between popular actors through their movie collaborations, using data from The Movie Database (TMDB) API. It constructs an actor graph and identifies the shortest path of relationships between any two actors.
Figure 2: UML Class Diagram for the Example Repository
Figure 3: UML Sequence Diagram for the Example Repository
@@ A.1.3 Architecture Design
.1 /. .2 .gitignore. .2 src. .3 acceptanceTest. .4 java. .5 Actor_relationship_game. .6 ARG_AcceptanceTest.java. .3 main. .4 java. .5 Actor_relationship_game. .6 Actor.java. .6 Movie.java. .6 ActorGraph.java. .6 ActorGraphUtil.java. .6 GameplayInterface.java. .6 GraphCreation.java. .6 TMDBApi.java. .3 test. .4 java. .5 Actor_relationship_game. .6 ActorTest.java. .6 MovieTest.java. .6 ActorGraphTest.java. .6 ActorGraphUtilTest.java. .6 GameplayInterfaceTest.java. .6 GraphCreationTest.java. .6 TMDBApiTest.java. .2 build.gradle. .2 README.md.
@@ A.3 Quality Assurance
Code Listing 1 is part of the unit test suite for the Actor Relationship Game Repository, specifically designed to validate the functionality of the Actor class. Using the JUnit framework, it defines two test cases: testActorIdAndName and testMovieIds. The first test, testActorIdAndName, instantiates an Actor object with a specific ID and name ("101" and "John Doe", respectively) and asserts that the getId() and getName() methods correctly return these values, ensuring the actor’s identity is accurately stored and retrievable. The second test, testMovieIds, creates another Actor object and adds two movie IDs ("201" and "202") to the actor’s list of movie IDs. It then verifies that these movie IDs are indeed associated with the actor by checking if the actor’s getMovieIds() set contains the added IDs. Together, these tests check the integrity of the Actor class’s basic functionalities: maintaining an actor’s identity and managing their associated movie IDs.
@@
In Code Listing 2, the example acceptance test is designed to verify the functionality of generating and comparing actor lists from graph data. It employs the runGradleTask method to execute specific Gradle tasks for creating graph data and generating actor lists into specified file paths, using parameters for file names to differentiate between reference and test data. The test first runs the runGraphCreation task with paths for both reference and test graphs, followed by the runActorGraphUtil task to generate actor lists from these graphs into specified file paths. Once the actor lists are generated, the test reads lines from both the reference and test actor list files and then iterates through each line in the reference actor list, followed by an assertion that each actor from the reference list is also present in the test list. This process effectively checks the integrity and consistency of the actor list generation feature by ensuring that the test actor list replicates the reference list accurately, thereby validating the application’s capability to process and output graph-related data correctly.
@@
Appendix B Dataset
B.1 Dataset Construction
B.2 Dataset Statistics
In Table 2, we present an exhaustive statistical breakdown of our datasets. DevEval contains a collection of 22 curated repositories, spanning across four widely-used programming languages (Python, C/C++, Java, JavaScript) and a diverse range of domains. The dataset is characterized by its multi-file structure. The Python repositories in our dataset are relatively straightforward, with each repository comprising approximately two files and an average of 276 lines of code. In contrast, the repositories pertaining to statically-typed programming languages, namely C/C++ and Java, are more complex, featuring an increased count of code files and lines. For JavaScript, our usage of the Vue.js framework necessitates that models adeptly navigate the framework’s templates and development paradigms. Consequently, JavaScript repositories exhibit the highest number of code files and lines, posing a substantial challenge for LLMs. Additionally, we have prepared extensive reference acceptance and unit tests for each repository to facilitate rigorous evaluation of the implementation task.
Appendix C The Baseline System
We introduce our baseline system formulated for DevEval, building upon the foundations of ChatDev Qian et al. (2023a, b). ChatDev is a virtual, chat-powered software development system that adheres to the conventional waterfall model. It bifurcates the development process into four primary tasks (dubbed as phrases in ChatDev): design, coding, testing, and documentation. Within this system, multiple LLM agents assume diverse roles such as programmers, reviewers, and testers, pertinent to each phase. ChatDev is characterized by its utilization of a chat chain mechanism, which segments each phase into smaller, atomic tasks. This approach enables context-sensitive, multi-turn dialogues between two distinct roles, facilitating the proposal and validation of solutions for individual tasks.
In contrast to ChatDev, our development of baseline system incorporates several features and enhancements. We have restructured the task design to align closely with the evaluation criteria of DevEval. Specifically, this includes the integration of comprehensive input to the system, exemplified by well-structured PRDs. This integration is crucial in addressing and examining the issue of hallucination in controlled experimental settings. Moreover, our baseline system expands upon the capabilities of ChatDev, supporting a wider range of tasks, including standard Object-Oriented programming designs (UML class and sequence diagrams), repository planning (architecture design), environment setup, and acceptance testing. A significant advancement in our baseline system is its compatibility with multiple programming languages and their corresponding runtime environments. This feature is coupled with the provision of comprehensive execution feedback to the system.
Implementation Details
We utilize LMDeploy for the deployment of CodeLlama and DeepSeek-Coder models. [^7] https://github.com/InternLM/lmdeploy Acknowledging the potential for extensive input context in DevEval tasks, we configure the context length to 32K for these models. For the Software Design task, we set the temperature parameter to 0.2, while for the remaining four tasks, we use a temperature of 0. Other hyperparameters in the experiment are maintained at default settings. All code-related tasks are rigorously evaluated in an isolated sandbox environment, utilizing Docker technology.
Appendix D Software Design Evaluation
We follow previous work Zheng et al. (2023a) to conduct a pairwise comparison to determine which response is better, focusing on the metrics of general principles and faithfulness (see the corresponding prompts in Figures 456). To reduce the expenditure of the OpenAI GPT API and human effort, the scope of our evaluation was confined to a subset of our dataset. This process involves 192 pairs across eight repositories, eight models, and three sub-tasks. Regarding the LLM judge, we use GPT-4-Turbo as the judge and GPT-3.5-Turbo as the baseline model. To mitigate the issue of position bias (Zheng et al., 2023a; Shi et al., 2024), i.e., LLM judges preferring response at a certain position regardless of the content, the evaluation was executed in a dual mode, evaluating each pair twice in different orders (384 pairs in total), with inconsistent decisions being considered as a tie. For the human evaluation, we shuffle the order of two responses and annotate each pair thrice to obtain the human majority.
Figure 4: Evaluation Guidance for UML Class.
Figure 5: Evaluation Guidance for UML Sequence.
Figure 6: Evaluation Guidance for Architecture Design.
The customized LLM-as-a-Judge prompt for evaluating software design is detailed in Fig. 7, structured to facilitate pairwise comparisons in accordance with predefined evaluation guidelines. The LLM’s judgments are extracted by employing regular expressions to identify the selection specified after "Choice:" within the judges’ responses.
Figure 7: LLM-as-a-Judge prompt for software design evaluation